# V-OCBF: Learning Safety Filters from Offline Data via Value-Guided Offline Control Barrier Functions

### v1

@@ 6 Conclusion, Limitations and Future works
A current limitation is that the framework is not explicitly trained to handle adversarial disturbances or worst-case model errors. Future work will explore robust extensions of V-OCBF that incorporate adversarial or uncertainty-aware training objectives, enabling safety guarantees under stronger disturbance models and broader distribution shifts.
@@ A.2.1 Why Expectile Regression in V-OCBF?
In V-OCBF, the barrier function ideally requires computing $\max_{u^{\prime}\in\mathcal{U}}B(x^{\prime})$, but performing this maximization directly is problematic in the offline setting because the dataset supports only a restricted action set $\mathcal{U}_{\mathcal{D}}=\{\,u\mid(x,u,x^{\prime})\in\mathcal{D}\,\}$, and querying $B(x^{\prime})$ for unseen actions $u^{\prime}\notin\mathcal{U}_{\mathcal{D}}$ introduces unsupported targets that can distort or destabilize learning. Moreover, even restricting the maximization to $\mathcal{U}_{\mathcal{D}}$ leads to a brittle update since a hard maximum is extremely sensitive to noise or rare outlier transitions, often producing poor barrier estimates. Expectile regression provides a smooth and data-supported alternative by emphasizing higher (safer) barrier values when $\tau>0.5$, effectively approximating the maximal barrier value over $\mathcal{U}_{\mathcal{D}}$ while never querying out-of-distribution actions, thereby yielding a stable and principled mechanism for synthesising offline CBVF.
@@ B.2.1 Qualitative Analysis
In continuation to the Hopper plot in the main paper (Fig. 4), we include plots for all the remaining MuJoCo safety gymnasium environments covered in this paper with the visualization of the episode rollouts, both with and without V-OCBF. Given the safety rates of V-OCBF compared to the other baselines (refer Fig. 3), it becomes important to visually realize the real-time effect on performance (Figures 6, 7, 8).
@@ C.2 Robustness to errors in Learned Dynamics
To evaluate the robustness of our pipeline to errors in the learned dynamics, we measure how additive perturbations affect the safety performance of the CBF-based controllers. Concretely, we perturb the computed safe action $\omega$ by adding zero-mean Gaussian noise $\mathcal{N}(0,\sigma^{2})$ with a range of standard deviations $\sigma$ (refer 5). For every noise level we run 500 independent trials (with 5 different random seeds each) and report the safety rate defined as the fraction of trials that remain within the safe set over the evaluation horizon. Table 5 summarizes the results across several CBF variants, for each method we report mean safety rate $\pm$ standard error. Across noise levels spanning small to large perturbations, the safety rates degrade only mildly, indicating that the learned CBFs tolerate moderate inaccuracies in the dynamics / control signal. These results suggest that, while precise dynamics models improve performance, our approach provides substantially safety-aware control even when the dynamics model is imperfect.
@@
C.2.1 Sensitivity to Expectile parameter ($\tau$)
@@
To investigate this effect, we conduct a study by comparing $\tau=0.5$ against higher expectile levels that emphasize the upper tail of the safety value distribution. Table 6 summarises the safety rates achieved on the AGV environment for different $\tau$ values:
@@
The results reveal two key patterns. First, the behavior-induced barrier ($\tau=0.5$) yields the lowest safety rate, consistent with its tendency to underrepresent safety-critical regions that are not frequently explored by the behavior policy. Second, as $\tau$ increases, the barrier becomes progressively more conservative in safety-critical areas and better aligned with the upper end of the demonstrated safety values, resulting in consistently higher safety rates.
The best performance is achieved near $\tau\rightarrow 1$, indicating that prioritizing the high-value portion of the demonstration distribution leads to safer barrier estimates and improved closed-loop performance.
Appendix D Experimental Details
D.1 Experimental Hardware
@@
D.2 Hyperparameters for the Proposed Algorithm
@@
D.3 Hyperparameters for the Baselines
For the CBF-based baselines (NCBF, iDBF, CCBF), we use the same network architectures and learned dynamics models described above to ensure a fair comparison. Hyperparameters for the remaining safe offline RL baselines (COptiDICE, BEAR-Lag, BC, FISOR) are provided in Table 8. We use the official implementations of BEAR-Lag and COptiDICE from liu2024dsrl and of FISOR from 1, and train all methods on the same DSRL datasets.
Table 8: Hyperparameters for all the Baselines network. Optimizer: Adam, LR: $1\times 10^{-3}$, Activation: ReLU (applies to all rows). For COptiDICE and BEAR-Lag refer OSRL, liu2024dsrl, for the official implementations. For FISOR refer 1 for official implementation.

### v2

@@ 6 Conclusion, Limitations and Future works
While V-OCBF reliably learns useful control-barrier candidates from offline demonstrations and enables QP-based safety filtering under unknown dynamics and limited action authority, the learned barrier is not a formally verified certificate, function approximation and learned-dynamics errors mean guarantees remain empirical. Furthermore, the framework is not explicitly trained to handle adversarial disturbances or worst-case model errors. To close this gap, future work will explore integration of formal-certification (e.g., Lipschitz-based verification (tayal2024learning) and conformal prediction (lindemann2025formal)). Moreover, we will explore robust extensions of V-OCBF that incorporate adversarial or uncertainty-aware training objectives, enabling safety guarantees under stronger disturbance models and broader distribution shifts.
@@ A.2.1 Why Expectile Regression in V-OCBF?
In V-OCBF, the barrier function ideally requires computing $\max_{u^{\prime}\in\mathcal{U}}B(x^{\prime})$, but performing this maximization directly is problematic in the offline setting because the dataset supports only a restricted action set $\mathcal{U}_{\mathcal{D}}=\{\,u\mid(x,u,x^{\prime})\in\mathcal{D}\,\}$, and querying $B(x^{\prime})$ for unseen actions $u^{\prime}\notin\mathcal{U}_{\mathcal{D}}$ introduces unsupported targets that can distort or destabilize learning. Moreover, even restricting the maximization to $\mathcal{U}_{\mathcal{D}}$ leads to a brittle update since a strict maximum is extremely sensitive to noise or rare outlier transitions, often producing poor barrier estimates. Expectile regression provides a smooth and data-supported alternative by emphasizing higher (safer) barrier values when $\tau>0.5$, effectively approximating the maximal barrier value over $\mathcal{U}_{\mathcal{D}}$ while never querying out-of-distribution actions, thereby yielding a stable and principled mechanism for synthesising offline CBVF.
@@
To illustrate the effect of the expectile parameter we show barrier functions learned for the AGV case study using $\tau\in\{0.5,0.6,0.7,0.8,0.9,0.99\}$ (Figure 4). As $\tau$ increases the volume of the sub-zero level set visibly shrinks: lower $\tau$ values produce larger sub-zero regions because the learning target is more influenced by poorer next-state outcomes present in the data, which inflates the set of states that appear unsafe under the learned barrier. By contrast, larger $\tau$ up-weights higher (safer) barrier targets within the dataset and yields a smoother, upper-envelope–like approximation to the maximization over dataset-supported actions; sampling actions from this higher expectile therefore produces a smaller sub-zero level set (i.e., a tighter unsafe region) and correspondingly larger estimated safe set.
Figure 4: $\mathbf{\tau}$-Expectiles: Figure shows a depiction of the learned Barrier Function for different $\tau$ values with Expectile Regression. With increasing value of $\tau~(>0.5)$, the feasible region increases, indicating a more accurate representation of the true barrier function.
A.3 Feasibility of V-OCBF-QP
Formal safety guarantees of standard QP-based CBF formulations generally hold under the assumption of unbounded actuation. In practice, however, in the presence of control bounds, a nominal CBF may conflict with feasibility requirements, causing the corresponding CBF-QP to become infeasible. To address this within our framework, we adopt a standard relaxation technique from the safe control literature by introducing a slack variable $\delta$. This relaxes the hard safety constraint when strictly necessary while heavily penalizing any violations. The resulting slack-augmented V-OCBF-QP is formulated as:
$\pi_{\mathrm{safe}}(x)$ $=\operatorname*{arg\,min}_{u\in\mathbb{U},\delta\in\mathbb{R}}\|u-\pi_{\text{ref}}(x)\|^{2}+\lambda\delta^{2}$ (22)
$\textrm{s.t. }\mathcal{L}_{f}B(x)+\mathcal{L}_{g}B(x)u+\kappa\left(B(x)\right)+\delta\geq 0,$
where $\lambda>0$ is a penalty coefficient.
Empirically, $\delta$ remains close to zero throughout our experiments. Because V-OCBF explicitly accounts for control constraints during learning, the synthesized policy naturally respects admissible bounds. Consequently, the safety filter rarely becomes infeasible, and any triggered corrective adjustments require minimal slack usage.
@@ B.2.1 Qualitative Analysis
In continuation to the Hopper plot in the main paper (Fig. 3), we include plots for all the remaining MuJoCo safety gymnasium environments covered in this paper with the visualization of the episode rollouts, both with and without V-OCBF. Given the safety rates of V-OCBF compared to the other baselines (refer Fig. 2), it becomes important to visually realize the real-time effect on performance (Figures 7, 8, 9).
@@ C.2 Robustness to errors in Learned Dynamics
To evaluate the robustness of our pipeline to errors in the learned dynamics, we measure how additive perturbations affect the safety performance of the CBF-based controllers. Concretely, we perturb the computed safe action $\omega$ by adding zero-mean Gaussian noise $\mathcal{N}(0,\sigma^{2})$ with a range of standard deviations $\sigma$ (refer Table 5). For every noise level we run 500 independent trials (with 5 different random seeds each) and report the safety rate defined as the fraction of trials that remain within the safe set over the evaluation horizon. Table 5 summarizes the results across several CBF variants, for each method we report mean safety rate $\pm$ standard error. Across noise levels spanning small to large perturbations, the safety rates degrade only mildly, indicating that the learned CBFs tolerate moderate inaccuracies in the dynamics / control signal. These results suggest that, while precise dynamics models improve performance, our approach provides substantially safety-aware control even when the dynamics model is imperfect.
@@
C.3 Sensitivity to Expectile parameter ($\tau$)
@@
To investigate this effect, we conduct a study by comparing $\tau=0.5$ against higher expectile levels that emphasize the upper tail of the safety value distribution. Table 6 summarises the safety rates achieved on the AGV environment for different $\tau$ values.
@@
The results reveal two key patterns. First, the behavior-induced barrier ($\tau=0.5$) yields the lowest safety rate, consistent with its tendency to underrepresent safety-critical regions that are not frequently explored by the behavior policy. Second, as $\tau$ increases, the barrier becomes progressively more conservative in safety-critical areas and better aligned with the upper end of the demonstrated safety values, resulting in consistently higher safety rates. Figure 4 depicts this nature of the learned barrier functions on varying the value of $\tau$.
Under idealized conditions (deterministic dynamics and no dataset disturbances ,i.e., when the best actions reliably lead to the best next states), pushing the expectile $\tau\to 1$ concentrates the estimator on the upper tail of the target distribution and, under favorable (nearly deterministic) conditions, can yield the strong empirical performance, prioritizing high-value next-state targets. However, this extreme setting is brittle to rare poor actions or stochastic disturbances that produce optimistic next states; in such cases an estimator that strictly targets the extreme upper tail may overfit those spurious outcomes.
To empirically validate this reasoning, we conducted an additional set of experiments evaluating the impact of $\tau$ across multiple Safety MuJoCo tasks using standard, public datasets from DSRL, which are not strictly deterministic. We evaluated both safety (measured via Cost) and task performance (measured via Reward) for $\tau\in\{0.5,0.7,0.9,0.99\}$, as shown in Table 7.
Table 7: Reward and Cost across different $\tau$ expectile values for selected MuJoCo environments
$\tau=0.5$ $\tau=0.7$ $\tau=0.9$ $\tau=0.99$
Environment Reward Cost Reward Cost Reward Cost Reward Cost
Ant 565.36 0.0 639.61 0.0 791.77 0.0 734.66 0.02
Half Cheetah 100.39 0.0 180.87 0.0 679.73 0.0 582.58 0.05
Hopper 96.94 0.0 126.54 0.0 769.69 0.0 526.72 0.27
As observed in Table 7, increasing $\tau$ toward $1$ (specifically $\tau=0.99$) results in degraded safety performance, reflected by increased Cost, and in some cases reduced Reward. This trend is consistent across the reported tasks and aligns with our theoretical intuition regarding sensitivity to stochastic transitions. In contrast, $\tau=0.9$ achieves a more favorable balance between reward maximization and safety preservation. Similar patterns were observed across other environments. Based on these findings, we selected $\tau=0.9$ for the experiments reported in the main paper.
C.4 Sensitivity to Dynamics Model Accuracy
To empirically evaluate the resilience of our framework against approximation errors in the learned dynamics, we conducted an analysis using the Autonomous Ground Vehicle environment to quantify how variations in model fidelity influence safety performance. We evaluated four distinct configurations representing varying levels of accuracy: the known system dynamics, the standard learned model used in our main experiments, and two degraded models trained for only 50% and 25% of the standard steps, respectively. We quantified the model quality by computing the mean squared prediction error over a set of randomly sampled state-action pairs, which confirmed a monotonic reduction in predictive accuracy across the models.
Table 8: Effect of Dynamics Model Accuracy on Safety Rates at the time of Inference (Mean over 500 episode rollouts).
Method Safe Episodes (%) Episode Reward Mean L2 Error
BC+V-OCBF (Known Dynamics) 99.52 $\pm$ 0.58 55.91 $\pm$ 0.39 0
BC+V-OCBF (Learned Dynamics) 98.28 $\pm$ 0.54 54.93 $\pm$ 0.46 $1.48\times 10^{-2}$
BC+V-OCBF (Half-Learned Dynamics) 97.03 $\pm$ 0.45 52.44 $\pm$ 0.18 $2.93\times 10^{-2}$
BC+V-OCBF (Quarter-Learned Dynamics) 96.32 $\pm$ 0.73 51.91 $\pm$ 0.32 $4.46\times 10^{-2}$
Subsequently, we utilized each model to synthesize the controller and evaluated the resulting empirical safety rates. As detailed in Table 8, the results indicate that the safety rate experiences only a modest reduction despite the increase in model prediction error. These findings suggest that V-OCBF maintains its effectiveness in mitigating safety violations even when the underlying dynamics model is imperfect, supporting the practical applicability of the method in settings where exact system identification is challenging.
Appendix D Additional Experiments on Boat Navigation
To further assess the generalization capabilities of V-OCBF, we evaluated the method on a boat navigation task characterized by nonlinear river drift and static obstacles. This environment, while low-dimensional, presents a control challenge due to the drift dynamics which can actively push the agent into unsafe regions if not properly anticipated. Such drift dynamics are particularly difficult for offline safe RL and CBF-based methods to handle when relying on learned models or indirect constraint formulations.
The system dynamics are given by:
$(x_{1})_{t+1}$ $=(x_{1})_{t}+(a_{1}+2-0.5x_{2}^{2})\cdot\Delta t$
$(x_{2})_{t+1}$ $=(x_{2})_{t}+a_{2}\cdot\Delta t$
where $\Delta t$ represents discrete time step, $a_{1},a_{2}$ represent the velocity control action in $x_{1}$ and $x_{2}$ directions respectively, subject to the actuation limit $a_{1}^{2}+a_{2}^{2}\leq 1$. The term $2-0.5x_{2}^{2}$ introduces a state-dependent drift along the $x_{1}$-axis. The safety constraints are encoded as:
$\ell(x):=min(\|x-(-0.5,0.5)^{T}\|-0.4,\|x-(-1.0,-1.2)^{T}\|-0.4)$ (29)
where $\ell(x)<0$ indicates that the boat is inside a obstacle, thereby ensuring that the sub-level set of $\ell(x)$ defines the failure region.
Table 9: Boat Navigation Experiment: Percentage Safe Episodes against additional baseline methods. Evaluated over 500 episodes and 5 seed values.
Method Safe Episodes (%)
BC+NCBF (Alexander2020NCBF) 88.43 $\pm$ 0.59
CPQ (xu2022constraints) 54.14 $\pm$ 0.26
WSAC (wei2024wsac) 79.34 $\pm$ 0.87
CDT (liu2023constrained) 69.94 $\pm$ 0.42
C2IQL (liu2025ciql) 70.77 $\pm$ 0.76
BEAR-Lag (kumar2019stabilizing) 71.69 $\pm$ 0.16
COptiDICE (lee2022coptidice) 87.88 $\pm$ 0.65
FISOR (1) 85.78 $\pm$ 0.22
BC+V-OCBF (Ours) 97.56 $\pm$ 0.13
We compared V-OCBF against a broader set of baselines, including CPQ (xu2022constraints), WSAC (wei2024wsac), CDT (liu2023constrained), and C2IQL (liu2025ciql), in addition to the methods evaluated in the main text. The results, summarized in Table 9, demonstrate that V-OCBF consistently achieves a higher empirical safety rate compared to the baseline methods. Notably, the strongest baseline exhibits a safety rate approximately 9% lower than our approach. These results, combined with the experiments presented in Section 5, indicate that V-OCBF provides a robust mechanism for minimizing safety violations across diverse dynamical systems, effectively handling the challenges posed by nonlinear drift and learned dynamics.
Appendix E Experimental Details
E.1 Experimental Hardware
@@
E.2 Hyperparameters for the Proposed Algorithm
@@
E.3 Hyperparameters for the Baselines
For all the CBF-based baselines (NCBF, iDBF, CCBF), we use the same network architectures and learned dynamics models described above to ensure a fair comparison. Hyperparameters for the safe offline RL baselines (COptiDICE, BEAR-Lag, BC, FISOR) are provided in Table 11. We use the official implementations of CPQ, CDT, BEAR-Lag and COptiDICE from (OSRL) liu2024dsrl, WSAC from wei2024wsac, C2IQL from liu2025ciql and FISOR from 1, and train all methods on the same DSRL datasets.
Table 11: Hyperparameters for all the Baselines network. For COptiDICE and BEAR-Lag refer OSRL, liu2024dsrl, for the official implementations. For FISOR refer 1 for official implementation.